# From Intent to Execution: Multimodal Chain-of-Thought Reinforcement Learning for Precise CAD Code Generation

### v1

@@ 1 Introduction
Computer-Aided Design (CAD) systems are foundational tools in engineering and manufacturing, enabling the creation of precise, parameterized 3D models for a wide range of industrial applications. However, traditional CAD workflows are highly manual, requiring extensive domain expertise and design iteration effort. While prior work has attempted to automate 3D reconstruction from sketches or images (Jayaraman et al. 2022; Li et al. 2020; Wu, Xiao, and Zheng 2021), the resulting models are typically non-editable, approximate, and lack parametric control—limitations that are incompatible with the tight tolerance requirements of industrial design. Recent advances in vision-language models (VLMs), particularly in natural language to code generation, open up a new paradigm: transforming high-level design intent into executable and semantically grounded CAD code. This capability promises to enable customized manufacturing, interpretable geometry modeling, and a substantial reduction in CAD development cost, while also bridging the gap between natural communication and formal design languages.
Despite this promising direction, translating complex human intent into executable CAD modeling code remains highly challenging. First, it necessitates understanding precise dimensions and physical constraints embedded in design language. Second, the model must generate code that is syntactically valid, structurally coherent, and functionally executable. Recent efforts begin to tackle this space. Text-Guided CAD-Coder(He et al. 2025) maps language prompts to pre-defined command sequences, but its rigid structure restricts generalization to a narrow set of shapes. VLM CAD-Coder(Doris et al. 2025) utilizes single-modality image input to generate CAD code, but the lack of accurate dimensional constraints leads to coarse approximations that fall short in real industrial practice. CAD-Coder (Guan et al. 2025) achieves high accuracy by generating structured CadQuery code from expert-written descriptions, but its reliance on expert-level inputs hinders accessibility for general users. Furthermore, all of these methods operate on single-modal input, limiting adaptability to diverse, real-world design scenarios.
@@ 2.1 RL Adaptation for Code-Oriented VLMs
Code generation from natural language is a central capability of modern LLMs (Jiang et al. 2024a). While supervised fine-tuning (SFT) based on maximum likelihood helps models learn syntax and structure, it does not directly optimize for downstream metrics such as execution correctness or semantic alignment. To address this, reinforcement learning (RL) is increasingly employed to align generation with task-specific objectives (Fujimoto, Meger, and Precup 2019). RL-based techniques for code LLMs generally fall into three categories: Policy gradient methods like CodeRL (Le et al. 2022) and LEDEX (Jiang et al. 2024b) use PPO (Schulman et al. 2017) to maximize execution-based rewards, but require extensive computation and careful reward design. Preference-based methods reframe RLHF as a binary classification task over response pairs. DPO (Rafailov et al. 2023) improves training efficiency via closed-form policy updates but struggles with nuanced preference modeling. Gradient-based reward optimization, such as DeepSeek-RL (Guo et al. 2024), scales to large models using in-batch advantage estimation. Hybrid approaches like Swe-RL (Wei et al. 2025) and StepCoder (Dou et al. 2024) combine supervised and reward-driven objectives to improve reasoning and correctness. Each method offers trade-offs between supervision cost, scalability, and alignment fidelity, making them complementary in different stages of model adaptation.
@@ 2.2 CAD Code Generation
CAD code generation aims to create precise, editable 3D models directly from user instructions. This approach improves on conventional 3D reconstruction by enabling parametric control and design flexibility. Several methods tackle this task using different modalities and representations. CAD-MLLM (Xu et al. 2024) and GenCAD (Alam and Ahmed 2024) map image inputs to CAD commands using vision-language models. Img2CAD (You et al. 2024) separates structure prediction and parameter regression in a two-stage design. CAD2Program (Wang et al. 2025) introduces a more flexible code format for richer modeling expressiveness. Recent methods increasingly adopt CadQuery, a Python-based modeling language, as the output format (He et al. 2025; Doris et al. 2025; Guan et al. 2025). Unlike command sequences or DSLs, CadQuery supports loops, conditionals, and parameterization, making it both expressive and executable. Its compatibility with industrial formats (e.g., STEP, STL) makes it a practical bridge between intent and deployable design.
@@ 3.1 Overview
Inspired by recent advancements in RL for code generation (Guo et al. 2024; Dou et al. 2024), we propose CAD-RL, a reinforcement learning post training framework tailored for executable, high-precision CAD code generation. CAD-RL takes flexible multimodal inputs, including both natural language descriptions (e.g., from non-expert users) and structured design language (e.g., from professional engineers), optionally augmented by reference images. This flexible input paradigm makes our system adaptable to diverse, real-world CAD design scenarios.
@@ 3.3 Reinforcement Learning Post Training
External Evaluation Reward $R_{\text{eval}}$: To complement the syntactic and geometric criteria captured by $R_{\text{exec}}$ and $R_{\text{geom}}$, we introduce an auxiliary reward that reflects high-level functional correctness and semantic fidelity from a task-centric perspective. This reward is derived from an external evaluator—specifically, GPT-4o (Achiam et al. 2023)—which provides a soft, interpretable score assessing how well the generated CAD code aligns with the design intent.
@@
Standard policy gradient methods, such as PPO (Schulman et al. 2017), constrain policy updates using a trust region to ensure training stability. This constraint prevents excessive divergence between the current policy $\pi_{\theta}$ and a reference policy $\pi_{\text{ref}}$, often initialized from a pretrained or SFT model.
@@
The original GRPO (Shao et al. 2024) framework computes loss at the sample level by first averaging token-level losses within each sequence, followed by aggregation across samples. In CAD practice, different tokens play markedly different roles in shaping geometric accuracy. Drawing errors such as numerical dimensions have disproportionately severe impacts on the correctness of the resulting 3D model. To address this, we introduce Precision Token Loss, a weighted loss function that amplifies gradients on semantically important tokens. Let $\omega_{t}$ denote the importance weight for token $s_{t}$, determined based on token type or positional heuristics. The loss over a batch $\mathcal{B}$ is computed as:
@@ 3.4 ExeCAD
LLaVA-1.5 (Liu et al. 2024) 0.2 178.19 110.76 1.67 0.0 144.76 91.30 0.13
Phi-3.5-Vision (Abdin et al. 2024) 4.4 192.09 121.05 32.50 6.7 161.9 93.45 51.63
InternVL2.5 (Chen et al. 2024) 13.7 161.75 86.56 55.17 19.5 136.29 72.13 72.33
Qwen2.5-VL (Team 2025) 30.6 103.70 1.95 52.47 31.6 114.64 7.70 62.30
Claude4 (Anthropic 2024) 38.9 92.81 4.96 63.26 41.8 91.29 4.37 64.42
InternVL3 (Zhu et al. 2025) 40.9 82.59 2.57 56.72 43.5 82.72 2.15 72.03
Gemini2.5 Pro (Google DeepMind 2024) 41.8 83.23 2.43 64.39 45.1 79.93 2.06 66.13
GPT-4o (Achiam et al. 2023) 47.3 77.28 1.76 73.16 47.7 65.52 1.47 72.72
CAD-Coder (Guan et al. 2025) $\ast$ 66.2 41.96 1.17 94.72 70.4 29.63 1.02 95.44
@@
We construct ExeCAD by refining two open-source datasets that are independently derived from the DeepCAD project: GenCAD-Code (Alam and Ahmed 2024), which provides syntactically valid but dimensionally noisy CadQuery scripts, and Text2CAD (Khan et al. 2024), which contains human-written natural language descriptions that are often vague, redundant, or semantically inconsistent. To ensure high-quality multimodal supervision, we employ GPT-4o as an alignment assistant and guide it through four refinement stages using prompt engineering:
@@ 4.3 Benchmarking CAD Code Generation
Table 1 presents a comprehensive evaluation on the ExeCAD benchmark across two types of textual inputs: Natural Language Description and Structured Design Language. The comparison includes eight state-of-the-art pretrained vision-language models (VLMs)—spanning both open-source models such as Qwen2.5-VL (Team 2025) and InternVL3 (Zhu et al. 2025), as well as proprietary models like GPT-4o (Achiam et al. 2023)—all evaluated without any task-specific post training. We also benchmark two fine-tuned baselines: CAD-Coder (Guan et al. 2025) and our proposed CAD-RL.
@@
4) Metric selection matters: Mean CD vs. Median CD vs. IoU. Following CAD-Coder (Guan et al. 2025), we adopt both the mean and median Chamfer Distance (CD) to evaluate shape similarity. While Mean CD is widely used, it is highly sensitive to outliers—especially when large parts are under-sampled due to limited point resolution, leading to disproportionate distortion of the average. Median CD, on the other hand, captures typical cases and offers a more robust estimation. However, it fails to reflect subtle but critical structural inaccuracies. For instance, Claude 4 (Anthropic 2024) reports a better IoU than Qwen2.5-VL (Team 2025) under natural language input, yet its median CD is significantly worse. To address this discrepancy, we incorporate IoU as an additional metric. Unlike perceptual or distributional distances, IoU directly enforces near-exact volumetric overlap, aligning more closely with industrial design requirements where precise geometric conformity is critical. Given the low tolerance for structural drift in CAD applications, IoU provides a more faithful measure of functional shape accuracy.

### v2

@@ 1 Introduction
Computer-Aided Design (CAD) systems are foundational tools in engineering and manufacturing, enabling the creation of precise, parameterized 3D models for a wide range of industrial applications. However, traditional CAD workflows are highly manual, requiring extensive domain expertise and design iteration effort. While prior work has attempted to automate 3D reconstruction from sketches or images (jayaraman2022solidgen; li2020sketch2cad; wu2021deepcad), the resulting models are typically non-editable, approximate, and lack parametric control—limitations that are incompatible with the tight tolerance requirements of industrial design. Recent advances in vision-language models (VLMs) (niu2025creft; niu2025pht), particularly in natural language to code generation, open up a new paradigm: transforming high-level design intent into executable and semantically grounded CAD code. This capability promises to enable customized manufacturing, interpretable geometry modeling, and a substantial reduction in CAD development cost, while also bridging the gap between natural communication and formal design languages.
Despite this promising direction, translating complex human intent into executable CAD modeling code remains highly challenging. First, it necessitates understanding precise dimensions and physical constraints embedded in design language. Second, the model must generate code that is syntactically valid, structurally coherent, and functionally executable. Recent efforts begin to tackle this space. Text-Guided CAD-Coder(he2025cad) maps language prompts to pre-defined command sequences, but its rigid structure restricts generalization to a narrow set of shapes. VLM CAD-Coder(doris2025cad) utilizes single-modality image input to generate CAD code, but the lack of accurate dimensional constraints leads to coarse approximations that fall short in real industrial practice. CAD-Coder (guan2025cad) achieves high accuracy by generating structured CadQuery code from expert-written descriptions, but its reliance on expert-level inputs hinders accessibility for general users. Furthermore, all of these methods operate on single-modal input, limiting adaptability to diverse, real-world design scenarios.
@@ 2.1 RL Adaptation for Code-Oriented VLMs
Code generation from natural language is a central capability of modern LLMs (jiang2024survey). While supervised fine-tuning (SFT) based on maximum likelihood helps models learn syntax and structure, it does not directly optimize for downstream metrics such as execution correctness or semantic alignment. To address this, reinforcement learning (RL) is increasingly employed to align generation with task-specific objectives (fujimoto2019off). RL-based techniques for code LLMs generally fall into three categories: Policy gradient methods like CodeRL (le2022coderl) and LEDEX (jiang2024ledex) use PPO (schulman2017proximal) to maximize execution-based rewards, but require extensive computation and careful reward design. Preference-based methods reframe RLHF as a binary classification task over response pairs. DPO (rafailov2023direct) improves training efficiency via closed-form policy updates but struggles with nuanced preference modeling. Gradient-based reward optimization, such as DeepSeek-RL (guo2024deepseek), scales to large models using in-batch advantage estimation. Hybrid approaches like Swe-RL (wei2025swe) and StepCoder (dou2024stepcoder) combine supervised and reward-driven objectives to improve reasoning and correctness. Each method offers trade-offs between supervision cost, scalability, and alignment fidelity, making them complementary in different stages of model adaptation.
@@ 2.2 CAD Code Generation
CAD code generation aims to create precise, editable 3D models directly from user instructions. This approach improves on conventional 3D reconstruction by enabling parametric control and design flexibility. Several methods tackle this task using different modalities and representations. CAD-MLLM (xu2024cad) and GenCAD (alam2024gencad) map image inputs to CAD commands using vision-language models. Img2CAD (you2024img2cad) separates structure prediction and parameter regression in a two-stage design. CAD2Program (wang20252d) introduces a more flexible code format for richer modeling expressiveness. Recent methods increasingly adopt CadQuery, a Python-based modeling language, as the output format (he2025cad; doris2025cad; guan2025cad). Unlike command sequences or DSLs, CadQuery supports loops, conditionals, and parameterization, making it both expressive and executable. Its compatibility with industrial formats (e.g., STEP, STL) makes it a practical bridge between intent and deployable design.
@@ 3.1 Overview
Inspired by recent advancements in RL for code generation (guo2024deepseek; dou2024stepcoder), we propose CAD-RL, a reinforcement learning post training framework tailored for executable, high-precision CAD code generation. CAD-RL takes flexible multimodal inputs, including both natural language descriptions (e.g., from non-expert users) and structured design language (e.g., from professional engineers), optionally augmented by reference images. This flexible input paradigm makes our system adaptable to diverse, real-world CAD design scenarios.
@@ 3.3 Reinforcement Learning Post Training
External Evaluation Reward $R_{\text{eval}}$: To complement the syntactic and geometric criteria captured by $R_{\text{exec}}$ and $R_{\text{geom}}$, we introduce an auxiliary reward that reflects high-level functional correctness and semantic fidelity from a task-centric perspective. This reward is derived from an external evaluator—specifically, GPT-4o (achiam2023gpt)—which provides a soft, interpretable score assessing how well the generated CAD code aligns with the design intent.
@@
Standard policy gradient methods, such as PPO (schulman2017proximal), constrain policy updates using a trust region to ensure training stability. This constraint prevents excessive divergence between the current policy $\pi_{\theta}$ and a reference policy $\pi_{\text{ref}}$, often initialized from a pretrained or SFT model.
@@
The original GRPO (shao2024deepseekmath) framework computes loss at the sample level by first averaging token-level losses within each sequence, followed by aggregation across samples. In CAD practice, different tokens play markedly different roles in shaping geometric accuracy. Drawing errors such as numerical dimensions have disproportionately severe impacts on the correctness of the resulting 3D model. To address this, we introduce Precision Token Loss, a weighted loss function that amplifies gradients on semantically important tokens. Let $\omega_{t}$ denote the importance weight for token $s_{t}$, determined based on token type or positional heuristics. The loss over a batch $\mathcal{B}$ is computed as:
@@ 3.4 ExeCAD
LLaVA-1.5 (liu2024improved) 0.2 178.19 110.76 1.67 0.0 144.76 91.30 0.13
Phi-3.5-Vision (abdin2024phi) 4.4 192.09 121.05 32.50 6.7 161.9 93.45 51.63
InternVL2.5 (chen2024internvl) 13.7 161.75 86.56 55.17 19.5 136.29 72.13 72.33
Qwen2.5-VL (qwen2.5-VL) 30.6 103.70 1.95 52.47 31.6 114.64 7.70 62.30
Claude4 (anthropic2024claude35) 38.9 92.81 4.96 63.26 41.8 91.29 4.37 64.42
InternVL3 (zhu2025internvl3exploringadvancedtraining) 40.9 82.59 2.57 56.72 43.5 82.72 2.15 72.03
Gemini2.5 Pro (gemini2.5pro2024) 41.8 83.23 2.43 64.39 45.1 79.93 2.06 66.13
GPT-4o (achiam2023gpt) 47.3 77.28 1.76 73.16 47.7 65.52 1.47 72.72
CAD-Coder (guan2025cad) $\ast$ 66.2 41.96 1.17 94.72 70.4 29.63 1.02 95.44
@@
We construct ExeCAD by refining two open-source datasets that are independently derived from the DeepCAD project: GenCAD-Code (alam2024gencad), which provides syntactically valid but dimensionally noisy CadQuery scripts, and Text2CAD (khan2024text2cad), which contains human-written natural language descriptions that are often vague, redundant, or semantically inconsistent. To ensure high-quality multimodal supervision, we employ GPT-4o as an alignment assistant and guide it through four refinement stages using prompt engineering:
@@ 4.3 Benchmarking CAD Code Generation
Table 1 presents a comprehensive evaluation on the ExeCAD benchmark across two types of textual inputs: Natural Language Description and Structured Design Language. The comparison includes eight state-of-the-art pretrained vision-language models (VLMs)—spanning both open-source models such as Qwen2.5-VL (qwen2.5-VL) and InternVL3 (zhu2025internvl3exploringadvancedtraining), as well as proprietary models like GPT-4o (achiam2023gpt)—all evaluated without any task-specific post training. We also benchmark two fine-tuned baselines: CAD-Coder (guan2025cad) and our proposed CAD-RL.
@@
4) Metric selection matters: Mean CD vs. Median CD vs. IoU. Following CAD-Coder (guan2025cad), we adopt both the mean and median Chamfer Distance (CD) to evaluate shape similarity. While Mean CD is widely used, it is highly sensitive to outliers—especially when large parts are under-sampled due to limited point resolution, leading to disproportionate distortion of the average. Median CD, on the other hand, captures typical cases and offers a more robust estimation. However, it fails to reflect subtle but critical structural inaccuracies. For instance, Claude 4 (anthropic2024claude35) reports a better IoU than Qwen2.5-VL (qwen2.5-VL) under natural language input, yet its median CD is significantly worse. To address this discrepancy, we incorporate IoU as an additional metric. Unlike perceptual or distributional distances, IoU directly enforces near-exact volumetric overlap, aligning more closely with industrial design requirements where precise geometric conformity is critical. Given the low tolerance for structural drift in CAD applications, IoU provides a more faithful measure of functional shape accuracy.